OpenStudy (anonymous):

A solid homogeneous cylinder of radius ‘r’ rolls without slipping on the inside of a stationary large cylinder of radius R. Find the equation of motion.

OpenStudy (anonymous):

Let the angular velocity of the system about the point os suspension at any time be ‘’ So, vc = (R – r) Again vc = r1 [where, 1 = rotational velocity of the sphere] 1 = r vc =        r R r  …(1) By Energy method, Total energy in SHM is constant. So, mg(R – r)(1 – cos) + (1/2) mvc 2+(1/2) I1 2 = constant  mg(R – r) (1 – cos) +(1/2) m(R – r)2 2 +(1/2) mr2 2 r R r     2 = constant  g(R – r) 1 – cos) + (R – r)2 2      5 1 2 1 = constant Taking derivative, g(R – r) sin  dt d  10 7 R – r)22 dt d   g sin  = 2 × 10 7 (R – r)  g sin  = 5 7 (R – r)   = 7(R r) 5gsin   = 7(R r) 5g      = 2 = 7(R r) 5g   = constant So the motion is S.H.M. Again  =  7(R r) 5g  T = 2 5g 7(R  r)